2. When a capacitor, charged to a p.d. of 400 V, is connected to a voltmeter having a resistance of 25 M-Ohm, the voltmeter reading is observed to have fallen to 50 V at the end of an interval of 2 minutes. Find the capacitance of the capacitor.  


1
Expert's answer
2020-07-20T15:13:32-0400

The capacitor is discharging through the resistor therefore the value of charge is meant to decrease over time.

Therefore, the capacitor discharge formula is should be used in this case

"V=V_0e^{\frac{-t}{RC}}"


Making RC subject of the formula


"RC=\frac{t}{ln\frac{V_0}{V}}"


"RC=\frac{120}{ln\frac {400}{50}}"


RC= 57.08 seconds(This is the time constant)


now divide the time constant by the resistance to get the value of capacitance


"C=\frac{57.08}{25\times 10^{6}}"


Therefore capacitance is "2.2832\times 10^{-6}" F